Find t to the nearest hundredth if $1786 becomes $2063 at 2.6% with interest compounded monthly.
Answer by stanbon(75887)

Find t to the nearest hundredth if $1786 becomes $2063 at 2.6% with interest compounded monthly.
----------------
A(t) = P(1+ (r/n))^(nt
2063 = 1786(1+0.026/12))^(12t)
1.55095 = (1.002167)^(12t)
12t log 1.002167 = log 1.5595
12t = (log1.5595)/(log1.002167)
12 t = 205.2822119
t = 17.1069
===============
Cheers,
Stan H.